# PKS 2131--021 – Discovery of Strong Coherent Sinusoidal Variations from Radio to Optical Frequencies: Compelling Evidence for a Blazar Supermassive Black Hole Binary

### v1

@@ Abstract
Haystack and Owens Valley Radio Observatory (OVRO) observations recently revealed strong sinusoidal total flux density variations that maintained coherence between 1975 and 2021 in the blazar PKS 2131--021 ($z=1.283)$. This was interpreted as possible evidence of a supermassive black hole binary (SMBHB). Extended observations through 2023 show coherence over 47.9 years, with an observed period $P_{\textrm{15 GHz}}=(1739.3\pm 1.2)\,{\rm days}$. We reject, with $p$-value = $5.3\times 10^{-7}$, the hypothesis that the variations are due to random fluctuations in the red noise tail of the power spectral density. There is clearly a constant-period physical phenomenon in PKS 2131--021 producing coherent intermittent sinusoidal flux density variations. We find the coherent sinusoidal intensity variations extend from below 2.7 GHz to optical frequencies, from which we derive an observed period $P_{\textrm{optical}}=(1764\pm 36)$ days. Across this broad frequency range there is a monotonic phase shift in the sinusoidal variations with frequency. The same coherent periodicity is possibly also observed at $\gamma$-ray energies. The importance of well-vetted SMBHB candidates to searches for gravitational waves is pointed out. We estimate the fraction of blazars that are SMBHB candidates to be $>1$ in 100. Thus monitoring programs covering tens of thousands of blazars could discover hundreds of SMBHB candidates.
@@ 1 INTRODUCTION
The 40 m Telescope of the Owens Valley Radio Observatory (OVRO) has been engaged in monitoring the total 15 GHz flux densities of $\sim 1830$ blazars since 2008 (Richards et al., 2011) on a 3–4 day cadence. The first statistically robust report of a strong supermassive black hole binary (SMBHB) candidate (PKS 2131--021) showing sinusoidal variations in its radio light curve was presented by O’Neill et al. (2022, hereafter Paper 1).
@@
SMBHBs are a direct result of hierarchical structure formation and an important ingredient for our understanding of the evolution of the Universe (e.g., Volonteri et al., 2003; Maiolino et al., 2023; Übler et al., 2024). In spite of the significant effort that has been put into finding SMBHBs (see D’Orazio & Charisi, 2023, for a recent review), the number of compelling candidates remains small. This is particularly true for blazars, active galactic nuclei (AGN) with jets oriented toward the observer on Earth (Blandford et al., 2019; Hovatta & Lindfors, 2019). Until recently, the most compelling blazar SMBHB candidate with a separation $\ll 1$ pc had been OJ 287, identified through more than 100 years of optical observations showing an outburst every $\sim$11 years due to the secondary black hole crossing the accretion disk of the primary (Sillanpaa et al., 1988; Valtonen et al., 2016; Dey et al., 2021). All-sky surveys in optical and $\gamma$-rays have produced more candidates (e.g., PG 1302--102, Graham et al., 2015; PG 1553+113, Ackermann et al., 2015), and future facilities like the Vera C. Rubin observatory are expected to produce many more (e.g., Xin & Haiman, 2021).
@@
In the radio regime, periodicity searches have been limited because of the scarcity of facilities capable of regularly observing a large number of blazars. The Owens Valley Radio Observatory 40-m Telescope blazar program (Richards et al., 2011) is currently the largest radio flux-density monitoring program. Given the quality and length of the OVRO light curves, we have engaged in a search for statistically significant periodic and quasi-periodic oscillations (QPOs). In Paper 1 we presented our first results on PKS 2131--021, providing evidence of a periodicity in this blazar. These results showed that the observed periodicity is highly unlikely to be due to random fluctuations and the red noise tail in the variability spectrum. In view of the stability of its period, PKS 2131--021 is a strong SMBHB candidate. In this paper, we present further evidence, which we find compelling, supporting the hypothesis of the SMBHB nature of PKS 2131--021.
@@
We analyze the radio to optical light curves of PKS 2131--021 in terms of a model due to Blandford (see Paper 1), which ascribes the sinusoidal variations in PKS 2131--021 to aberration caused by the orbital motion in an SMBHB of the SMBH producing the jet. This same model had been developed by Sobacchi et al. (2017) in the context of a “spine-sheath” model, but we missed the underlying similarity. We will refer to this as the “Kinematic Orbital” model (KO model).
@@
As discussed in Paper 1, blazars showing evidence of periodicities should not be considered bona fide periodic or quasi-periodic oscillator (QPO) candidates unless it can be shown that there is a very low global probability ($p$-value $<1.3\times 10^{-3}$), taking into account the look-elsewhere effect, that the periodicity is due to red noise tail fluctuations, through a careful analysis of the power spectral density (PSD) and through simulations that reproduce both the observed PSD and the observed probability density function (PDF) seen in the light curve. Such analysis is essential for assessing the chance occurrence probability of sinusoidal features in light curves of AGN.
@@
We also analyze the light curves of PKS 2131--021 at 2.7 GHz, 7.9 GHz, 15.5 GHz and 31.4 GHz using Haystack observations from 1975–1983 (O’Dea et al., 1986), 15 GHz observations from OVRO, and ALMA observations at 91.5 GHz, 104 GHz, and 345 GHz. In addition, we analyze infrared data from WISE, optical data from the Zwicky Transient Facility (ZTF), and $\gamma$-ray data from the Fermi Gamma-ray Space Telescope, and show that sinusoidal variations, coherent with the radio variations, are clearly seen at optical wavelengths, and, furthermore, hints of coherent periodic variations are seen at both infrared wavelengths and $\gamma$-ray energies.
@@
Figure 1: Verified prediction in the 15 GHz light curve of PKS 2131--021. (a) Black points: The OVRO 15 GHz light curve presented in Paper 1, which extends up to the vertical dashed line; blue curve: the least-squares sine-wave fit to the Haystack+OVRO data given in Table 1 of Paper 1, extrapolated to the end of 2023. Note the divergence of the data (red points) from the fit. The prediction at this point was that the light curve would revert to the sine wave. The more recent 2023 observations (green points) show that the observed curve has returned to the extrapolated fit (blue curve), as predicted in Paper 1. The purple dashed sine curve, which is almost indistinguishable from the blue sine curve, is the best least squares fit, given in Table 1, to the Haystack data plus all of the OVRO data to the end of 2023. (b) The residual 15 GHz light curve after subtraction of the sinusoidal component (purple dashed sine wave) plotted in (a). The black horizontal line indicates the mean flux density in the residual.
@@
This paper is organized as follows: in Section 2 we discuss the continuation of the 15 GHz sinusoidal variations in PKS 2131--021 for another 2 yr; in Section 3 we discuss the radio spectrum, structure, and polarization of PKS 2131--021 to provide background for the later sections; in Section 4 we present evidence that strong coherent sinusoidal variations are seen from 2.7 GHz to optical frequencies, with a monotonically varying phase shift as a function of frequency; in Section 5 we discuss the X-ray observations of PKS 2131--021 and the prospects for future X-ray observations; in Section 6 we discuss the $\gamma$-ray emission from PKS 2131--021; in Section 7 we discuss the multifrequency observations in terms of the KO model; in Section 8 we discuss the fraction of blazars that are SMBHB candidates; in Section 9 we summarize our findings and discuss their implications. There are three appendices: in Appendix A we discuss the Doppler factor for the compact radio component; in Appendix B we discuss the implications of the absence of higher harmonics in the light curve; in Appendix C we discuss the determination of the parameters of the sinusoidal optical light curves.
@@ 2 The Continuation of the 15 GHz Sine Wave Signal in PKS 2131--021
The OVRO 15 GHz light curve of PKS 2131--021 is shown in Fig. 1, where the vertical dashed line marks the extent of our calibrated data (up to MJD 59197) analyzed in Paper 1. The sine-wave fitting method is described in Paper 1. In Fig. 1(a) the blue curve shows the least squares sine wave fit to the data prior to MJD 59197 (black points) extrapolated forward to the end of 2023 (MJD 60310). As can be seen, by mid-January 2023 (MJD 59959) the new data appeared at first to be diverging from the extrapolated sine curve (red points), but they returned to it in 2023 (green points), thus fulfilling the prediction, from Paper 1, of a constant long-term period. The observed 47.9 year continuity, when combined with the Haystack data, is unprecedented in the light curves of blazars. In Fig. 1(b) we show the residual variations in PKS 2131-021 after subtraction of the sinusoidal component.
The combination of the Haystack and OVRO 15 GHz data is shown in Fig. 2, together with two sine-wave fits. Between these two epochs the sinusoidal variation disappeared (see Paper 1). In Fig. 2(a) we show a sine-wave fit in which each data point is given equal weight. While the fit is quite good, it is clear that there is a slight offset in phase between the sine wave and the Haystack data. The Haystack data are particularly relevant with respect to any phase offsets because of the long time interval of 25 years that separates them from the OVRO data. For the purpose of determining the underlying periodicity, and hence the degree of coherence in the combined data set, in Fig. 2(b) we have assigned a weight of 879/154 = 5.71 to each of the Haystack observations, corresponding to the number of data points in the Haystack data (154) and the OVRO data (879). The weights of the OVRO data points are all 1. As can be seen in Fig. 2(b), this has brought the sine wave into alignment with the Haystack data, with no discernible difference in the fit to the OVRO data. The corresponding period is $P=(1739.2\pm 1.2)$ days, which may be compared with the period of $P=(1737.9\pm 2.6)$ days obtained in Paper 1. The period in the $z=1.283$ rest frame of the host galaxy is $P_{0}=762\pm 0.5$ days. As discussed in Paper 1, the cause of the difference in the amplitudes of the sine waves in the Haystack data and the OVRO data is unknown.
We have used the above best estimate of the period, $P=(1739.2\pm 1.2)$ days, as a prior in determining the best fitting sine wave to the OVRO data. The resulting parameters are given in Table 1, and the corresponding sine wave is shown by the purple dashed sine wave in Fig. 1(a).
Figure 2: Sine wave fits (red curves) to the combined Haystack observations from 1975 to 1983 (O’Dea et al., 1986) and OVRO observations from 2008 to the end of 2023. The amplitudes have been permitted to vary between the two data sets, but the period is derived from the combined data set. (a) Fit to the unweighted data. (b) Fit in which the Haystack observations have been assigned weights of $w=879/154=5.71$, where 879 and 154 are the numbers of data points in the OVRO and Haystack light curves, respectively (see text).
Table 1: The Best Least Squares OVRO Sine Wave
@@
Table 2: Probabilities and Significance Levels of GLS Tests Computed from Simulations with Matched Red Noise Tail
Test Test GLS $\mathcal{P}_{\rm peak}$ Period Range ($\Delta P$) Total Number of simulations p-value Significance
Number (max=1) (days) Simulations that pass test ($\sigma$)
1 OVRO data 0.81 All 300000 40 $1.33\times 10^{-4}$ 3.65
2 Haystack data $\geq 0.50$ 1657.5 – 1854.1 10000 197 $1.97\times 10^{-2}$ 2.06
3 phase - - - - 0.2 $0.8$
- Combining 1, 2 and 3 - - - - $5.25\times 10^{-7}$ $4.88$

### v2

@@ Abstract
Haystack and Owens Valley Radio Observatory (OVRO) observations recently revealed strong, intermittent, sinusoidal total flux-density variations that maintained coherence between 1975 and 2021 in the blazar PKS 2131--021 ($z=1.283)$. This was interpreted as possible evidence of a supermassive black hole binary (SMBHB). Extended observations through 2023 show coherence over 47.9 years, with an observed period $P_{\textrm{15 GHz}}=(1739.8\pm 17.4)\,{\rm days}$. We reject, with $p$-value = $2.09\times 10^{-7}$, the hypothesis that the variations are due to random fluctuations in the red noise tail of the power spectral density. There is clearly a physical phenomenon in PKS 2131--021 producing coherent sinusoidal flux density variations. We find the coherent sinusoidal intensity variations extend from below 2.7 GHz to optical frequencies, from which we derive an observed period $P_{\textrm{optical}}=(1764\pm 36)$ days. Across this broad frequency range there is a smoothly-varying monotonic phase shift in the sinusoidal variations with frequency. Hints of periodic variations are also observed at $\gamma$-ray energies. The importance of well-vetted SMBHB candidates to searches for gravitational waves is pointed out. We estimate the fraction of blazars that are SMBHB candidates to be $>1$ in 100. Thus monitoring programs covering tens of thousands of blazars could discover hundreds of SMBHB candidates.
@@ 1 INTRODUCTION
The 40 m Telescope of the Owens Valley Radio Observatory (OVRO) has been dedicated to monitoring the total 15 GHz flux densities of $\sim 1830$ blazars since 2008 (Richards et al., 2011) on a 3–4 day cadence. The first statistically robust report of a strong supermassive black hole binary (SMBHB) candidate (PKS 2131--021) showing sinusoidal variations in its radio light curve was presented by O’Neill et al. (2022, hereafter Paper 1).
@@
SMBHBs are a direct result of hierarchical structure formation and an important ingredient for our understanding of the evolution of the Universe (e.g., Volonteri et al., 2003; Maiolino et al., 2023; Übler et al., 2024). In spite of the significant effort that has been put into finding SMBHBs (see D’Orazio & Charisi, 2023, for a recent review), the number of compelling candidates remains small. This is particularly true for blazars, active galactic nuclei (AGN) with jets oriented toward the observer on Earth (Blandford et al., 2019; Hovatta & Lindfors, 2019). Until recently, the most compelling blazar SMBHB candidate with a separation $\ll 1$ pc had been OJ 287, identified through more than 100 years of optical observations showing an outburst every $\sim$11 years due to the secondary black hole crossing the accretion disk of the primary (Sillanpaa et al., 1988; Valtonen et al., 2016; Dey et al., 2021). All-sky surveys in optical and $\gamma$-rays have produced more candidates (e.g., PG 1302--102, Graham et al., 2015; PG 1553+113, Ackermann et al., 2015; Abdollahi et al., 2024; MAGIC Collaboration et al., 2024), and future facilities like the Vera C. Rubin observatory are expected to produce many more (e.g., Xin & Haiman, 2021).
@@
In the radio regime, periodicity searches have been limited because of the scarcity of facilities capable of regularly observing a large number of blazars. The OVRO 40-m Telescope blazar program (Richards et al., 2011) is currently the largest radio flux-density monitoring program. Taking advantage of the quality and length of the OVRO light curves, we have undertaken a search for statistically significant periodic and quasi-periodic oscillations (QPOs). In Paper 1 we presented the first results on PKS 2131--021, demonstrating highly statistically significant periodicity in this blazar. These results showed that the observed periodicity is unlikely to be due to random fluctuations and the red noise tail in the variability spectrum. In view of the stability of its period, PKS 2131--021 is a strong SMBHB candidate. In this paper, we present further evidence, which we find compelling, supporting the hypothesis of the SMBHB nature of PKS 2131--021.
@@
We analyze the radio to optical light curves of PKS 2131--021 in terms of a model due to Blandford (see Paper 1), which ascribes the sinusoidal variations in PKS 2131--021 to aberration caused by the orbital motion in an SMBHB of the black hole producing the jet. This same model had been developed by Sobacchi et al. (2017) in the context of a “spine-sheath” model, but we missed the underlying similarity. We will refer to this as the “Kinematic Orbital” model (KO model).
@@
As discussed in Paper 1, blazars showing evidence of periodicities should not be considered bona fide periodic or QPO candidates unless it can be shown that there is a very low global probability ($p$-value $<1.3\times 10^{-3}$), taking into account the look-elsewhere effect, that the periodicity is due to red noise tail fluctuations, through a careful analysis of the power spectral density (PSD) and through simulations that reproduce both the observed PSD and the observed probability density function (PDF) seen in the light curve. Such analysis is essential for assessing the chance occurrence probability of sinusoidal features.
@@
We also analyze the multi-frequency light curves of PKS 2131--021 using Haystack observations from 1975–1983 (O’Dea et al., 1986) at 2.7 GHz, 7.9 GHz, 15.5 GHz and 31.4 GHz; 15 GHz observations from OVRO; and ALMA observations at 91.5 GHz, 104 GHz, and 345 GHz. In addition, we analyze infrared data from WISE, optical data from the Zwicky Transient Facility (ZTF), X-ray data from Swift-XRT, and $\gamma$-ray data from the Fermi Gamma-ray Space Telescope, and show that sinusoidal variations, coherent with the radio variations, are clearly seen at optical wavelengths, and, furthermore, hints of coherent periodic variations are seen at both infrared wavelengths and $\gamma$-ray energies.
@@
Figure 1: Verified prediction in the OVRO 15 GHz light curve of PKS 2131--021. (a) Black points: The OVRO 15 GHz light curve presented in Paper 1, which extends up to the vertical dashed line; blue curve: the least-squares sine-wave fit to the Haystack+OVRO data given in Table 1 of Paper 1, extrapolated to the end of 2023. Note the divergence of the data (red points) from the fit. The prediction at this point was that the light curve would revert to the sine wave. The more recent 2023 observations (green points) show that the observed curve has returned to the extrapolated fit (blue curve), as predicted in Paper 1. The purple dashed sine curve, which is almost indistinguishable from the blue sine curve, is the best least squares fit, given in Table 1, to the Haystack data plus all of the OVRO data to the end of 2023. (b) The residual 15 GHz light curve after subtraction of the sinusoidal component (purple dashed sine wave) plotted in (a). The black horizontal line indicates the mean flux density in the residual.
A critical aspect of the sinusoidal variations in PKS 2131--021 reported in Paper 1 is the fact that they disappeared completely for 19 years (1984–2003). This raises two questions: (i) given this disappearance, how can we be sure that there is an underlying “clock” in PKS 2131--021, and (ii) what is responsible for the disappearance and re-appearance of the sinusoidal variations? The answer to (i) is that when the sinusoidal variations returned after the 19-year gap they had both the same period and were “in step” with the original sine wave. This indicates that the “clock” had continued ticking throughout. The rigorous statistical analysis of Paper 1 showed that the hypothesis that the two sets of sinusoidal variations (from 1975 to 1983 and from 2003 to 2021) arose from random fluctuations with a red noise power spectral density (PSD) could be rejected, with a $p$-value$=1.58\times 10^{-6}$ or a significance level of $4.66\sigma$. One possible answer to (ii), which was proposed in Paper 1, is that the sporadic nature of the observed sinusoidal variations could be due to variations in the fueling supplied to the region of the jet inside the compact core component that is producing the sinusoidal variations.
@@
This paper is organized as follows: in Section 2 we discuss the continuation of the 15 GHz sinusoidal variations in PKS 2131--021 for another 2 yr; in Section 3 we discuss the effects of correlated noise on our analysis; in Section 4 we discuss the statistical significance of our results; in Section 5 we discuss the radio spectrum, structure, and polarization to provide background for the later sections; in Section 6 we present evidence that strong coherent sinusoidal variations are seen from 2.7 GHz to optical frequencies, with a smoothly-varying monotonic phase shift as a function of frequency; in Section 7 we discuss the X-ray observations and the prospects for future X-ray observations; in Section 8 we discuss the $\gamma$-ray emission; in Section 9 we discuss the multifrequency observations in terms of the KO model; in Section 10 we discuss the fraction of blazars that are SMBHB candidates; in Section 11 we summarize our findings and discuss their implications. There are four appendices: in Appendix A we discuss the determination of the parameters of the sinusoidal optical light curves; in Appendix B we discuss the effects of non-white noise deviations from the sinusoidal features in the light curves; in Appendix C we discuss the Doppler factor for the compact radio component; in Appendix D we discuss the implications of the absence of higher harmonics in the light curve.
@@ 2 The Continuation of the 15 GHz Sine Wave Signal in PKS 2131--021
The OVRO 15 GHz light curve of PKS 2131--021 is shown in Fig. 1, where the vertical dashed line marks the extent of our calibrated data (up to MJD 59197) analyzed in Paper 1.
In Fig. 1(a) the blue curve shows a least squares sine wave fit (see section 2.1) to the data prior to MJD 59197 (black points) extrapolated forward to the end of 2023 (MJD 60310). As can be seen, by mid-January 2023 (MJD 59959) the new data appeared at first to be diverging from the extrapolated sine curve (red points), but they returned to it in 2023 (green points), thus fulfilling the prediction, from Paper 1, of a constant long-term period. The observed 47.9 year continuity, when combined with the Haystack data, is unprecedented in the light curves of blazars. In Fig. 1(b) we show the residual variations in PKS 2131--021 after subtraction of the sinusoidal component.
It is important to note that the significant departures from the sinusoidal variation in Fig. 1(a) in 2014, 2016 and 2022 do not undermine the SMBHB interpretation of this object. These departures are most clearly shown in the light curve with the sine wave subtracted in Fig. 1(b). There is nothing unusual in the blazar light curve of Fig. 1(b) (see e.g. Medd et al., 1972). It is mildly varying, and we have many such examples in the $\sim$1830 blazars in the 40 m Telescope monitoring sample. Thus it is to be expected that blazar SMBHBs will show deviations from pure sinusoidal variations due to the non-sinusoidally varying components, some of which are seen clearly as coming from different components in the very long baseline interferometry (VLBI) maps. This point is addressed in detail in Paper 1.
Figure 2: Sine wave fits (red curves) to the combined Haystack observations from 1975 to 1983 (O’Dea et al., 1986) and OVRO observations from 2008 to the end of 2023. The amplitudes have been permitted to vary between the two data sets, but the period is derived from the combined data set. (a) Fit to the unweighted data. (b) Fit to the weighted data (see text).
Table 1: The Least Squares OVRO Sine Wave
@@
2.1 The OVRO+Haystack period fit
We fitted the following functional model to the data:
$S_{{\rm model},i}=A\sin(\phi_{i}-\phi_{0})+S_{0}+n_{i},$ (1)
where $\phi_{i}=2\pi(t_{i}-t_{0})/P$ is the phase of the $i$th data point, $\phi_{0}$ is the phase at reference time $t_{0}$, $A$ is the amplitude, $S_{0}$ is the flux density offset, and $n_{i}$ is the noise contribution to the $i$th data point. We allow separate values for $A$ and $S_{0}$ for the OVRO and Haystack sections of the data. As discussed in Paper 1, the cause of the difference in the amplitudes of the sine waves in the Haystack data and the OVRO data is unknown, but, given that we ascribe the disappearance of the sinusoidal fluctuations from 1984 to 2003 to a drop in fueling, we suggest that the different amplitudes in these two epochs could be due to different levels of fueling.
The noise term $n_{i}$ includes observational noise and source intrinsic variability that is not part of the sinusoidal variations. The non-sinusoidal components are effectively a type of correlated noise as far as the fit is concerned, and this correlation should ideally be taken into account in the fit. In the following we apply two noise models. One treats the noise terms as uncorrelated, the other models the correlation. We will show that the fitted parameters are consistent, but the corresponding uncertainties depend on the choice of the noise model.
2.2 White Noise Model Fit
First, we treated all noise terms as independent Gaussian noise, $n_{i}\sim\mathbb{N}(0,\sigma_{i}^{2}+\sigma_{0}^{2})$, where $\sigma_{i}$ are the data uncertainties and $\sigma_{0}$ is the source intrinsic noise level, which may be different for OVRO and Haystack fits. The combination of the Haystack and OVRO 15 GHz data is shown in Fig. 2, together with two sine wave fits. Between these two epochs the sinusoidal variation disappeared (see Paper 1). In Fig. 2 (a) we show a sine-wave fit in which each data point is given equal weight. While the fit is quite good, it is clear that there is a slight offset in phase between the sine wave and the Haystack data. The Haystack data are particularly relevant with respect to any phase offsets because of the long time interval of 25 years that separates them from the OVRO data. For the purpose of determining the underlying periodicity, and hence the degree of coherence in the combined data set, in Fig. 2 (b) we have assigned a weight of 850/154 = 5.5 to each of the Haystack observations, corresponding to the number of data points in the Haystack data (154) and the OVRO data (850). The weights of the OVRO data points are all 1. We refer to this fit as a “weighted fit”. This has brought the sine wave into alignment with the Haystack data, with no discernible difference in the fit to the OVRO data. The corresponding period is $P=(1739.2\pm 1.2)$ days, which may be compared with the period of $P=(1737.9\pm 2.6)$ days obtained in Paper 1. The period for the weighted fit in the $z=1.283$ rest frame of the host galaxy is $P_{0}\!=\!P/(1+z)\!=\!761.8\pm 0.5$ days.
Table 1 shows the parameters of the best fit to the OVRO data alone, where we fixed the period at $P=(1739.2\pm 1.2)$ days, as estimated above from the weighted OVRO+Haystack fit.
3 The Effects of Correlated Noise
In Fig. 3 we show the light curves of PKS 2131--021 from 2.7 GHz to optical frequencies. In this section we discuss the light curves from centimeter to sub-millimeter wavelengths shown in Figs. 3(a) and (b). We discuss the optical light curves of 3(c) in §6.2. In Fig.3(a) we show the Haystack multifrequency light curves from O’Dea et al. (1986). We downloaded ALMA data from the ALMA calibrator website. These are shown in Fig. 3(b). The color-coded curves are least-squares fits of sine waves to the data, derived by the procedure presented in Paper 1 and in Appendix B.
We now consider the effects of correlated noise in these multi-frequency light curves. Our adopted noise model is based on light curve simulations that are further explained in Sec. 4. Details of the model fit are given in Appendix B. With this model, we get a maximum-posterior period of $P=(1739.8\pm 17.4)$ days. As explained in Appendix B the uncertainty is likely to be overestimated. Note that this estimate has almost exactly the same central value as the weighted white noise fit of $P=(1739.2\pm 1.2)$ of the previous section, showing that the period we measure is robust, and demonstrating that, when the correlation of the non-sinusoidal variations is taken into account, no higher weighting of the Haystack data is required in order to bring the phases into alignment across the 19-year gap in the sinusoidal variations.
Note that the white noise model fit estimates error bars that are lower by about an order of magnitude than those of the correlated noise fit. An analysis carried out by Hincks et al. (in preparation) on sub-millimeter blazar light curves, shows that the uncertainties provided by the sine wave fitting approach are reasonably accurate in the presence of Type-A variations, but better knowledge is needed of SMBHB phenomenology to determine the uncertainties to better than a factor of two or three.
3.1 The Correlated Noise Phase Shift with Frequency
In Fig. 3(a) and (b) we clearly see non-sinusoidal random variations in the signal on all time-scales. Furthermore, we see clear examples of non-sinusoidal variations that are correlated across different observing frequencies (see the dashed lines in Fig. 3b).
When measuring the phase differences of the sinusoidal variations between different frequency bands, it matters how these variations correlate between frequencies. We consider two cases: Type-A variations are correlated across frequency with the same time delay as the sinusoidal signal. Such variations could be caused, for example, by variations in fueling. Since these behave like the signal itself, they are not random as far as the phase fit is concerned. In a sinusoidal fit, Type-A variations enter as noise in the phase measurements, but this noise would be strongly correlated between bands, and would mostly cancel. The cancellation would be imperfect if the light curves have only partial overlap when measuring relative offsets. Type-B variations are not correlated between frequencies, or correlate with a different phase shift to that of the sinusoidal signal, as might happen due to a shock traveling obliquely across the jet. These contribute directly to the error bars of the relative phase measurements.
Due to the paucity of multi-wavelength light curves of SMBHB blazars over periods of several decades, we cannot reliably distinguish between Type-A and Type-B variations, so it is not possible to carry out a full analysis of the Type-A and Type-B variations, and their effect on the derived sine wave period and its uncertainty. A correlated noise fit as in Appendix B would need to be based on how the correlated noise at different frequencies correlate with each other. A fit based on the correlated noise for each light curve in isolation greatly overestimates the uncertainty of the relative phase offset between frequencies.
If the correlated non-sinusoidal variations are dominated by Type-A errors, as appears to be the case, then we can either:
1. Fit the lightcurve at each frequency as a scaled and offset version of the OVRO light curve, or
2. Fit a phase per light curve using a sinusoidal model with the period fixed at the value for the best fitting OVRO light curve during the period of the overlapping multi-frequency light curves, assuming white noise. It is important to consider only the period of overlap because of the changes in period with time due to long-term random variations, as described in detail in Paper 1, and as can be seen clearly in Fig. 1(b).
[…]
Figure 3: PKS 2131--021 light curves at radio, mm, sub-mm, and optical wavelengths showing a steady monotonic progression of the phase of the sinusoidal variations with frequency. (a) Haystack 2.7–31.4 GHz observations (O’Dea et al., 1986) together with the least-squares sine-wave fits to the data at each of the frequencies. There is a steady progression in phase with frequency between the four light curves, with the low-frequency sine waves lagging the high-frequency sine waves. (b) OVRO 15 GHz light curve plus ALMA light curves at 91.5, 104, and 345 GHz. As in (a), there is a continuing progression in the phases of the sine wave fits with frequency. The light gray dashed lines connect several Type-A correlated noise variations across various frequency bands (see text). (c) Comparison of the ZTF optical and OVRO radio light curves of PKS 2131--021. The ZTF $i$, $r$, and $g$-band light curves of PKS 2131--021 are shown. The large symbols are average values for each year. The black points show the OVRO 15 GHz points over this period, which provide the phase reference point. The curves show the least squares sine wave fits to the corresponding data. We associate the optical peak with the nearest radio peak, not the preceding radio peak. (d) The observed phase shifts, relative to the 15.5 GHz (Haystack) and 15 GHz (OVRO) light curves, of the sine-wave fits to the data, given in Table 2. A positive phase shift indicates that the light curve is shifted to a later time than the 15 GHz light curve. The curve shows a quadratic polynomial fit to the phase offset vs. log(frequency) data (see §6.4).
Table 2: Phase Shift with Frequency Relative to the 15.5 GHz and 15 GHz Light Curves
Frequency Phase Uncertainty
Band (Fraction (Fraction
of a cycle) of a cycle)
2.7 GHz 0.145 0.028
7.9 GHz 0.048 0.013
15 and 15.5 GHz 0 0
31.4 GHz $-0.034$ 0.022
91.5 GHz $-0.073$ 0.005
104 GHz $-0.077$ 0.005
345 GHz $-0.134$ 0.009
infrared $-0.27$ 0.05
Optical $-0.352$ 0.008